# Efficient Biological Data Acquisition through Inference Set Design

### v2

@@ 4.1 Visualizing Inference Set Design
To develop intuition for inference set design we first run experiments on the MNIST dataset. The whole MNIST training set is used as an inference set from which agents can acquire samples. The MNIST test set is split 50-50 into a validation set used for early stopping and a test set used for measuring model performance on held-out data inaccessible by agents. In many real-world applications, data used for training ML models contains samples that are inherently difficult, only partially observed, or mislabeled. We investigate these three cases and simulate the challenge of partial observability by removing the lower part of the image (Figure 2b) and simulate labeling noise by shuffling labels for selected digits (see Figure 2c).
@@
In these experiments we study the performance of an active agent using the least-confidence (LC) acquisition function of Equation 3 and a random agent, selecting samples with uniform probability over the inference set. At every active learning step an agent acquires a batch of 1000 samples from the inference set, adds them to the training set, and retrains the model from scratch. During the active learning experiment we monitor the accuracy on the inference set, test set, acquired batch, and the system accuracy as a whole (see Equation 1). We also record which samples have been acquired at every acquisition steps. In all three studied settings in Figure 2, we observe the same phenomenon: the generalisation accuracy measured on the held-out test set quickly saturates for both agents, but the active agent is able to obtain significant improvements on the inference set which are also reflected in individual acquisition batches and on the system’s overall performance.
@@
The mechanism at play here stems from the active design of the inference set by the active agent. We illustrate this for the case of noisy labeling function in Figure 4. By selectively acquiring the images it tends to mislabel, the active agent makes predictions on remaining samples in the inference set much easier. Since the majority of samples acquired by the active agent in the first 20 steps represent these challenging examples, its accuracy on the acquired batch is significantly lower than a random agent’s (Figure 2c, $3$rd row). However, once the process of removing mislabeled samples from the inference set is complete, the active agent’s accuracy on the acquired batch drastically increases and surpasses the random agent. To further explore the tradeoff between task difficulty and efficiency gains, we run experiments with shuffled labels for increasingly larger subsets of digits (see Figure 3). The results show that with a higher number of such “noisy” samples, the gap between the active and random agents on the inference set increases. The active agent is able to achieve 100% accuracy on the inference set even with shuffled labels for 80% of the data. However, a more corrupted target set also necessitates more acquisition batches to prune out the difficult examples from the inference set, incurring smaller experimental budget reductions for highly challenging settings.
4.2 Assessing Robustness of the Mechanism
@@
In the previous section, we have shown that inference set design design leads to significantly improved system performance compared to random sampling, even with highly corrupted datasets. In this section, we validate that these observations hold for a wildly different type of data – molecular datasets. We use the Quantum Machine 9 (QM9) (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014). QM9 contains $134k$ small organic molecules and their quantum chemical properties computed with Density Functional Theory (DFT). For inputs, we convert the SMILES strings molecular representations into their Extended Connectivity Fingerprints (ECFPs). We predict the HOMO-LUMO gap framed as a classification task by discetizing the values into two balanced classes using median as a boundary condition to explore agents’ performance on classification task. Our main active agent uses the least-confidence acquisition function. We also explore the use of the BALD acquisition function (Gal et al., 2017). In addition to active and random agents, for both datasets, we also evaluate the performance of heuristic-based acquisition orders (original sorting, smallest to largest molecules, sorted by SA-score, etc.) as well as diversity sampling (selecting molecules with largest Tanimoto distance to the acquired set).
@@ 4.3 Inference Set Design in the Real World
Modern HTS platforms combined with genetic perturbation techniques facilitate large scale cell microscopy experiments that are designed to capture the effects of biological perturbations (Celik et al., 2022). In these experiments before taking microscopy images, cells undergo perturbation by either a CRISPR/Cas9-mediated gene knock-out or injection of a bioactive molecular compound at given concentration. The obtained cell microscopy images are then processed and passed through a neural network to obtain embeddings that correspond to a specific perturbation. With data from such HTS pipeline it is possible to build “Maps of Biology” that contain organized information about known and new biological relationships (see Figure 7b). Each point on the map corresponds to the cosine similarity between perturbation embeddings and reveals how strongly two perturbations are related. Expanding maps of biology might uncover new biological relationships that in turn could guide the discovery of leads for new medicines. However, even with ultra-HTS platforms, experimentally acquiring microscopy images of all possible cell perturbations is unfeasible. We aim to apply inference set design to this problem with the goal of reducing costs associated with building maps of biology.
@@
Figure 8: Performance on pheno-similarity classification task on our proprietary phenomics dataset. The least-confidence agent is shown to be the most effective on this task and reaches the targeted system accuracy of 98% at around 80% of the acquisition procedure.
For our experiments, we start by using the publicly available RxRx3 dataset (Fay et al., 2023) which contains the learned embeddings for $17,063$ CRISPR/Cas9-mediated gene knock-out perturbations, as well as $1,674$ FDA-approved compound perturbations at $8$ concentrations each. We train our models to predict the cosine similarity between the perturbation embeddings of any gene-compound pair, and frame this as a classification problem. For each gene-compound pairs, we select the compound-concentration perturbation that yields the largest cosine-similarity with the target gene. The cosine-similarities are discretized by computing a gene-specific threshold using the upper semi-interquartile range (SIQR) with a step-size of $2.5$. Finally, we remove genes that have activity below 25%. The result is a $151\times 1,674$ matrix of labels encoding either pheno-similarity ($y=1$) or an absence of relationship ($y=0$) between a gene and a compound perturbation (see Figure 7b).
For this prediction task, we treat all gene embeddings as known and focus only on the much larger space of compound acquisition. To featurize compounds we use embeddings of MolGPS model (Sypetkowski et al., 2024). At every active learning step, agents acquire a batch of compounds from the inference set and uncover their relationships with all genes in the dataset (entire row of the matrix). The results on RxRx3 dataset show only minor improvements over random agent (see Appendix E). We believe this is due to high complexity of biological relationships and limited size of the inference set. Indeed, inferring cellular similarity between gene and compound perturbations from only a few hundred examples is highly unlikely. To push this analysis further, we rerun the same experiment but on our proprietary phenomics dataset. It is similar to RxRx3 but much larger. It contains cell perturbation embeddings for 102,855 compound and 5,580 gene perturbations, leading a $2,000$ times larger relationship matrix. Using inference set design, the active agent is again able to outperform the baselines on this task, as shown in Figure 8. These experiments represent promising results for large scale applications of active learning to hybrid screens for early drug discovery.
@@ 4.4 Empirical validation of weak calibration and stopping criterion
Figure 9: Model calibration evaluation on the inference set in our QM9 experiment for the least-confidence acquisition function. The models tend to be underconfident early in the acquisition loop, and overconfident later on, but remains weakly calibrated (monotonic ordering). At stopping time $t=\tau$, the model is both highly confident and highly accurate on the remaining samples in the inference set.
In this last experimental section, we validate the theoretical claims made in Section 2.2. The proposed stopping criterion assumes weak calibration i.e. that the true accuracy of the model does not need to be exactly equal to its confidence (perfect calibration), but that the ordering is conserved, i.e. that the true accuracy is a monotonically increasing function of the model’s confidence. Figure 9 shows that this assumption is empirically verified in our QM9 experiments and Figure 13 in Appendix F.1 confirms the same observation in our phenomics and MNIST experiments, supporting the claim that weak calibration can be achieved in practice. We also empirically validate (see Figure 14, Appendix F.2) that the proposed stopping criterion making use of the bound presented in Equation 5 does indeed cause the agent to stop acquiring new labels at a system accuracy superior to the targeted threshold (see Equation 4).
@@ 5 Discussion
While active learning has undergone significant theoretical development and experimentation over the past 30 years, its widespread adoption in the industry remains a challenge (Reker, 2019). Active learning is often framed as a means to improve generalization performance on held-out test sets, which is a goal made difficult by real-world challenges such as noise in labeling functions, high diversity, and partial observability of the input space. However, we believe that a shift in perspective can help unlock the full potential of active learning. By focusing on optimizing performance on a target set of samples and allowing the model to decide which examples to label and which to predict, active learning can become a more powerful tool for real-world applications.
@@ Acknowledgments
We are extremely grateful to the entire team at Recursion and Valence Labs for participating in insightful discussions and providing feedback on this work. We also wish to specifically thank Austin Tripp for offering valuable input on this work as well as Oscar Mendez-Lucio and Peter McLean for their precious support with data processing and for engaging in fruitful exchanges about predictive models for phenomics.
@@ A.1 Proof of Lemma 1
and the selected batch consists of the first $N_{b}$ samples. The remaining inference set is composed of the remaining $N-N_{b}$ samples, i.e. $\mathcal{D}^{t}_{\text{inf}}:=\mathcal{D}^{t-1}_{\text{inf}}\setminus\mathcal{D}^{t}_{b}=\{(x_{i},y_{i})\}_{i=1+N_{b}}^{N}$.
@@
$\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}P_{\mathcal{D}_{\text{inf}}^{t-1}}[Y=\hat{Y}|\hat{v}_{i}]\leq\frac{1}{N-N_{b}}\sum_{i=1+N_{b}}^{N}P_{\mathcal{D}_{\text{inf}}^{t-1}}[Y=\hat{Y}|\hat{v}_{i}]$ (13)
By definition, $\mathcal{D}_{\text{inf}}^{t}:=\mathcal{D}_{\text{inf}}^{t-1}\setminus\mathcal{D}_{b}^{t}$, the left-hand side from Equation 13 thus captures all the terms from $\mathcal{D}_{b}^{t}$ and the right-hand side those from $\mathcal{D}_{\text{inf}}^{t}$. Rewriting each side using Equation 11, we obtain:
@@ C.1 Molecular dataset preprocessing
In many practical applications exact geometries of screened molecules are unknown as they require computationally expensive DFT calculations. As a first data processing step, we use RDKit library (Landrum et al., 2024) to convert molecular structures into SMILES strings and compute their Extended Connectivity Fingerprints (ECFPs). The SMILES representation provides complete information about molecule’s composition and atomic connectivity, however, it removes all information about 3D atomic positions. Using SMILES representation is a common solution that simplifies generation of candidate molecules for screening but makes property prediction a more challenging task as many properties vary depending on specific 3D conformation of a molecule.
Both molecular datasets are cleaned by removing duplicated SMILES and fingerprints as well as single-atom structures. For total energies in Molecules3D dataset we use reference correction technique where atomic energies are calculated using linear model fitted to the counts of atoms in a molecule of each element present in the dataset (obtained atomic energies are presented in Table 1). For reference correction a randomly selected sample of 100k molecules is used. The atomic energies are then subtracted from the total energies of all molecules in the dataset. The obtained referenced-corrected energies are normally distributed with mean around 0 eV. A small number of outliers with reference-corrected energy values above 10 standard deviations are removed from the dataset as well as 100k samples that were used for reference correction to avoid data leakage.
The final QM9 and Molecule3D datasets contain $133,885$ and $3,453,538$ molecules respectively. Both datasets are split into inference, validation, and test sets with 80%, 5%, 15% fractions. The QM9 HOMO-LUMO gap values are discretized into 2 balanced classes using median as a boundary condition to explore agents’ performance on classification task.
@@ C.2 RxRx3 dataset preprocessing
The RxRx3 dataset contains one embedding for each well. Each perturbation type (gene-guide pair or compound-concentration pair) has several replicates across wells, plates and experiments. Each plate also contains unperturbed control cells which are used to keep track of and eliminate a portion of the batch effects (Sypetkowski et al., 2023). These raw embeddings thus need to be aligned and aggregated. We align them by centering and scaling each perturbation embedding to the embeddings of the experiment-level unperturbed control wells. The embeddings are then aggregated through a multi-stage averaging procedure, across wells, plates, experiments and guides (for CRISPR perturbations), which yields an average embedding for each gene-perturbation and each compound-concentration perturbation. We then use the obtained embeddings to compute cosine similarities between gene and compound perturbations in the RxRx3 dataset.
@@ Appendix D Hyperparameters and Implementation details
For all presented experiments in this work we use MLP models with residual connections (Touvron et al., 2021). All experiments were repeated with 3 different random seeds. Hyperparameters for each experiments are summarized in Table 2.
@@
Number of ensemble members None None 5 None None
@@ E.1 Regression Task on Molecules3D
To evaluate the inference set design paradigm on a regression task we use a large dataset Molecules3D (Xu et al., 2021). Molecules3D contains structures and DFT-computed properties of approximately $4$ million molecules. In our experiments we aim to predict the HOMO-LUMO gap and total energy. For inputs, we convert the SMILES strings molecular representations into their Extended Connectivity Fingerprints (ECFPs).
@@
Although, in many applications predicting properties of larger molecules present a more challenging task, our experiments on the Molecules3D dataset demonstrate that acquiring molecules ordered from large to small may harm the predictions on inference set and overall system performance (see Figure 10). One of the reasons is the distribution of chemical elements across molecules in the dataset. When acquiring molecules from small to large, all unique chemical elements of the Molecules3D dataset are present in the training set after acquiring just the first 1,000 samples. However, when acquiring molecules from large to small, some chemical elements remain only in the inference set until the very end of the experiment which is especially detrimental for the total energy predictions (see Figure 11). This result demonstrates that using heuristic rules such as ordering molecules by size for data acquisition does not guarantee optimal acquisition or generalization of heuristic rules to new data.

### v3

@@ 4.1 Visualizing Inference Set Design
To develop intuition for inference set design we first run experiments on the MNIST dataset. The whole MNIST training set is used as the target set from which agents can acquire samples. The MNIST test set is split 50-50 into a validation set used for early stopping and a test set used for measuring model performance on held-out data inaccessible by agents. In many real-world applications, data used for training ML models contains samples that are inherently difficult, only partially observed, or mislabeled. We investigate these three cases. Specifically, we simulate the challenge of partial observability by removing the lower part of the image (Figure 2b) and simulate labeling noise by shuffling labels for selected digits (see Figure 2c).
@@
In these experiments we study the performance of an active agent using the least-confidence (LC) acquisition function of Equation 3 and a random agent, selecting samples with uniform probability over the inference set. At every active learning step an agent acquires a batch of $1000$ samples from the inference set, adds them to the training set, and retrains the model from scratch. Across these active learning steps, we monitor the accuracy on the inference set, the test set, the acquired batch, and the system accuracy as a whole (see Equation 1). We also record which samples have been acquired at every acquisition step. In all three studied settings in Figure 2, we observe the same phenomenon: the generalisation accuracy measured on the held-out test set quickly saturates for both agents, but the active agent is able to obtain significant improvements on the inference set, and these gains are also reflected in individual acquisition batches and on the system’s overall performance.
@@
The mechanism at play here stems from the active design of the inference set by the agent. We illustrate this for the case of a noisy labeling function in Figure 4. By selectively acquiring the images it tends to mislabel, the active agent makes predicting the labels on the remaining samples in the inference set much easier. Since the majority of samples acquired by the active agent in the first 20 steps represent these challenging examples, its accuracy on the acquired batch is significantly lower than a random agent’s (Figure 2c, $3$rd row). However, once the process of removing mislabeled samples from the inference set is complete, the active agent’s accuracy on the acquired batch drastically increases and surpasses the random agent. To further explore the tradeoff between task difficulty and efficiency gains, we run experiments with shuffled labels for increasingly larger subsets of digits (see Figure 3). The results show that with a higher number of such “noisy” samples, the gap between the active and random agents on the inference set increases. The active agent is able to achieve 100% accuracy on the inference set even with shuffled labels for 80% of the data. However, a more corrupted target set also necessitates more acquisition batches to prune out the difficult examples from the inference set, incurring smaller experimental budget reductions for highly challenging settings.
4.2 Assessing the Robustness of the Mechanism
@@
In the previous section, we have shown that inference set design leads to significantly improved system performance compared to random sampling, even with highly corrupted datasets. In this section, we validate that these observations hold for a wildly different type of data – molecular datasets. We use the Quantum Machine 9 (QM9) (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014). QM9 contains $134k$ small organic molecules and their quantum chemical properties computed with Density Functional Theory (DFT). For inputs, we convert the SMILES strings molecular representations into their Extended Connectivity Fingerprints (ECFPs). We predict the HOMO-LUMO gap framed as a classification task by discretizing the values into two balanced classes using median as a boundary condition to explore agents’ performance on classification task. Our main active agent uses the least-confidence acquisition function. We also explore the use of the BALD acquisition function (Gal et al., 2017). In addition to active and random agents, for both datasets, we also evaluate the performance of heuristic-based acquisition orderings (original ordering, smallest to largest molecules, ordered by SA-score, etc.) as well as a diversity sampling strategy (selecting molecules with the largest Tanimoto distance to the acquired set).
@@ 4.3 Inference Set Design in the Real World
Modern HTS platforms combined with genetic perturbation techniques facilitate large scale cell microscopy experiments that are designed to capture the effects of biological perturbations (Celik et al., 2022). In these experiments before taking microscopy images, cells undergo perturbation by either a CRISPR/Cas9-mediated gene knock-out or injection of a bioactive molecular compound at a given concentration. The obtained cell microscopy images are then processed and passed through a neural network to obtain embeddings that correspond to a specific perturbation. With data from such HTS pipeline it is possible to build “Maps of Biology” that contain organized information about known and new biological relationships. Each point on the map corresponds to the cosine similarity between perturbation embeddings and encodes whether two perturbations are related. Expanding maps of biology might uncover new biological relationships that in turn could guide the discovery of leads for new medicines. However, even with ultra-HTS platforms, experimentally acquiring microscopy images of all possible cell perturbations is unfeasible. We aim to apply inference set design to this problem with the goal of reducing costs associated with building maps of biology.
@@
Figure 8: Performance on a pheno-similarity classification task on our proprietary phenomics dataset. The least-confidence agent is shown to be the most effective on this task and reaches the targeted system accuracy of 98% at around 80% of the acquisition procedure.
For our experiments, we start by using the publicly available RxRx3 dataset (Fay et al., 2023) which contains the learned embeddings for $17,063$ CRISPR/Cas9-mediated gene knock-out perturbations, as well as $1,674$ FDA-approved compound perturbations at $8$ concentrations each. We frame this as a classification problem and train our models to predict the discretized cosine similarity between the perturbation embeddings of any gene-compound pair. As preprocessing, for each gene-compound pairs, we select the compound-concentration perturbation that yields the largest cosine-similarity with the target gene. The cosine-similarities are discretized by computing a gene-specific threshold using the upper semi-interquartile range (SIQR) with a step-size of $2.5$. Finally, we remove the genes that have an activity rate (across compounds) below $25\%$. The result is a $151\times 1,674$ matrix of labels encoding either pheno-similarity ($y=1$) or an absence of relationship ($y=0$) between a gene and a compound perturbation (see Figure 7b).
For this prediction task, we treat all gene embeddings as known and focus only on the much larger space of compound acquisition. To featurize compounds we use embeddings of the MolGPS model (Sypetkowski et al., 2024). At every active learning step, agents acquire a batch of compounds from the inference set and uncover their relationships with all genes in the dataset (entire row of the matrix). The results on the RxRx3 dataset show only minor improvements over a random agent (see Appendix E). We believe this is due to high complexity of biological relationships and limited size of the inference set. Indeed, successfully learning to infer cellular similarity between gene and compound perturbations from only a few hundred examples is highly unlikely. To push this analysis further, we re-run the same experiment but on our proprietary phenomics dataset. It is similar to RxRx3 but much larger. It contains cell perturbation embeddings for 102,855 compound and 5,580 gene perturbations, leading to a $2,000$ times larger relationship matrix. Using inference set design, the active agent is now able to outperform the baselines on this task, as shown in Figure 8. These experiments represent promising results for large scale applications of active learning to hybrid screens for early drug discovery.
@@ 4.4 Empirical validation of weak calibration and stopping criterion
In this last experimental section, we validate the theoretical claims made in Section 2.2. The proposed stopping criterion assumes weak calibration i.e. that the true accuracy of the model does not need to be exactly equal to its confidence (perfect calibration), but that the correct ordering is conserved. Figure 12 in Appendix F.1 shows that this assumption is empirically verified in all of our experiments, supporting the claim that weak calibration can be achieved in practice. In Figure 13, in Appendix F.2, we also empirically validate that the proposed stopping criterion making use of the bound presented in Equation 5 does indeed cause the agent to stop acquiring new labels at a system accuracy superior to the targeted threshold (see Equation 4).
@@ 5 Discussion
While active learning has undergone significant theoretical development and experimentation over the past 30 years, its widespread adoption in the industry remains a challenge (Reker, 2019). Active learning is often framed as a means to improve generalization performance on held-out test sets, which is a goal made difficult by real-world challenges such as noisy labeling functions as well as high diversity and partial observability of the input space. However, we believe that a shift in perspective can help unlock the untapped potential of active learning. By focusing on optimizing performance on a target set of samples and allowing the model to decide which examples to label and which to predict, active learning can become a powerful tool for real-world applications.
@@ Acknowledgments
We are grateful to the entire team at Recursion and Valence Labs for participating in insightful discussions and providing feedback on this work. We also wish to specifically thank Austin Tripp for offering valuable input on this work as well as Oscar Mendez-Lucio and Peter McLean for their precious support with data processing and for engaging in fruitful exchanges about predictive models for phenomics.
@@ A.1 Proof of Lemma 1
and the selected batch consists of the first $N_{b}$ samples. The remaining inference set is composed of the remaining $N-N_{b}$ samples, i.e. $\mathcal{D}^{t}_{\text{inf}}:=\mathcal{D}^{t-1}_{\text{inf}}\setminus\mathcal{D}^{t}_{b}=\{(x_{i},y_{i})\}_{i=N_{b}+1}^{N}$.
@@
$\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}P_{\mathcal{D}_{\text{inf}}^{t-1}}[Y=\hat{Y}|\hat{v}_{i}]\leq\frac{1}{N-N_{b}}\sum_{i=N_{b}+1}^{N}P_{\mathcal{D}_{\text{inf}}^{t-1}}[Y=\hat{Y}|\hat{v}_{i}]$ (13)
By definition of $\mathcal{D}_{\text{inf}}^{t}:=\mathcal{D}_{\text{inf}}^{t-1}\setminus\mathcal{D}_{b}^{t}$, the left-hand side from Equation 13 thus captures all the terms from $\mathcal{D}_{b}^{t}$ and the right-hand side those from $\mathcal{D}_{\text{inf}}^{t}$. Rewriting each side using Equation 11, we obtain:
@@ C.1 Molecular dataset preprocessing
In many practical applications, exact geometries of screened molecules are unknown as they require computationally expensive DFT calculations. As a first data processing step, we use the RDKit (Landrum et al., 2024) and Molfeat (Noutahi et al., 2023) libraries to convert molecular structures into SMILES strings and compute their Extended Connectivity Fingerprints (ECFPs). Both molecular datasets are cleaned by removing duplicated SMILES and fingerprints as well as single-atom structures. For total energies in the Molecules3D dataset we use a reference correction technique where atomic energies are calculated using a linear model fitted to the counts of atoms of each type present in a molecule (the obtained atomic energies are presented in Table 1). For reference correction, a randomly selected sample of 100k molecules is used. The atomic energies are then subtracted from the total energies of all molecules in the dataset. The obtained referenced-corrected energies are normally distributed with mean around 0 eV. A small number of outliers with reference-corrected energy values above 10 standard deviations are removed from the dataset as well as the 100k samples that were used for reference correction to avoid data leakage.
The final QM9 and Molecule3D datasets contain $133,885$ and $3,453,538$ molecules respectively. Both datasets are split into inference, validation, and test sets with 80%, 5%, 15% fractions. The QM9 HOMO-LUMO gap values are discretized into 2 balanced classes using median as a boundary condition to explore agents’ performance on a classification task.
@@ C.2 RxRx3 dataset preprocessing
The RxRx3 dataset contains one embedding for each well. Each perturbation type (gene-guide pair or compound-concentration pair) has several replicates across wells, plates and experiments. Each plate also contains unperturbed control cells which are used to keep track of and eliminate a portion of the batch effects (Sypetkowski et al., 2023). These raw embeddings thus need to be aligned and aggregated. We align them by centering and scaling each perturbation embedding to the embeddings of the experiment-level unperturbed control wells. The embeddings are then aggregated through a multi-stage averaging procedure, across wells, plates, experiments and guides (for CRISPR perturbations), which yields an average embedding for each gene-perturbation and each compound-concentration perturbation. We then use the obtained embeddings to compute the cosine similarities between gene and compound perturbations in the RxRx3 dataset.
@@ Appendix D Hyperparameters and Implementation details
For all experiments presented in this work we use MLP models with residual connections (Touvron et al., 2021). All experiments were repeated with 3 different random seeds and the hyperparameters are summarized in Table 2.
@@
Number of ensemble members - - 5 - -
@@ E.1 Regression Task on Molecules3D
To evaluate the inference set design paradigm on a regression task we use the Molecules3D dataset (Xu et al., 2021). Molecules3D contains structures and DFT-computed properties of nearly $4$ million molecules. In our experiments we aim to predict the HOMO-LUMO gap and total energy. For inputs, we convert the SMILES strings molecular representations into their Extended Connectivity Fingerprints (ECFPs) (see Appendix C.1 for all details).
@@
Although, in many applications predicting properties of larger molecules presents a more challenging task, our experiments on the Molecules3D dataset demonstrate that acquiring molecules ordered from large to small may harm the predictions on inference set and overall system performance (see Figure 9). One of the reasons is the distribution of chemical elements across molecules in the dataset. When acquiring molecules from small to large, all unique chemical elements of the Molecules3D dataset are present in the training set after acquiring just the first 1,000 samples. However, when acquiring molecules from large to small, some chemical elements remain only in the inference set until the very end of the experiment which is especially detrimental to the total energy predictions (see Figure 10). This result demonstrates that using heuristic rules such as ordering molecules by size for data acquisition does not guarantee optimal acquisition or generalization of heuristic rules to new data and instead risks harming the performance due to the introduction of such biases.